OpenStudy (geometry_hater):

Find the number of sides for a regular polygon if the area is 480, the side length is 10 and the apothem is 6. Find the number of sides for a regular polygon if the area is 480, the side length is 10 and the apothem is 6. @Mathematics

OpenStudy (anonymous):

The area can be computed using only the number of sides and the side length. A regular polygon with 8 sides of length 10 has an area of\[200 \text{ Cot}\left[\frac{\pi }{8}\right] \]or 482.8 to the nearest tenth of a square unit. The area for a polygon of 7 sides and 9 sides, side length of 10, is 363.4 and 618.2 square units respectively.